OpenStudy (anonymous):

[3.06] Given the function f(x) = 2(x + 4), find x if f(x) = 20.

OpenStudy (anonymous):

Let 20 = 2(x + 4) > 20 = 2x + 8 {expansion} > Rearrange to find x > 2x = 20 - 8 > x = 12 / 2